OpenStudy (anonymous):

Solve the polynomial inequality x^4>13x^2-36 and express the answer in interval notation

OpenStudy (lgbasallote):

let a = x^2 a^2 = x^4 a^2 > 13a - 36 a^2 - 13a + 36 > 0 (a-9)(a-4) > 0 the roots would be a = 9 and a =4 therefore a <9 and a>4 now can you do the rest and put it in interval notation?

OpenStudy (anonymous):

thats where i was left of at....idk wat to do from then on

OpenStudy (lgbasallote):

here's a hint \[2<x<3\] if i put this in interval notation \[(2,3)\] that means it's between 2 and 3 but 2 and 3 are not included because i used parenthesis. if 2 and 3 are included then it should be [2,3 ] does that help?